OpenStudy (anonymous):

Select the equation of the line that passes through the point (3, -1) and is parallel to the line on the graph.

OpenStudy (anonymous):

Wheres the equations? o-o

OpenStudy (anonymous):

We also need to see the graph lol

OpenStudy (anonymous):

im new to this so i dont know how to copy and paste images but its a horizontal line between 2 and 3 so about 2.5

OpenStudy (anonymous):

You can attach a file, o-o im new too. but i figured everything out xD yay me.

OpenStudy (anonymous):

did it ^-^

OpenStudy (anonymous):

Wheres the Equations? ;o just copy and paste them.

OpenStudy (anonymous):

Question 5 (Multiple Choice Worth 1 points) (03.07) Graph with a line going through points zero comma two point five and four comma two point five. Select the equation of the line that passes through the point (3, -1) and is parallel to the line on the graph. y = -1 y = 3 y = x -1 y = 3x - 1

OpenStudy (texaschic101):

the slope of a horizontal line is 0. A parallel line will have the same slope. y = mx + b slope(m) = 0 (3,-1)...x = 3 and y = -1 now we sub -1 = 0(3) + b -1 = b and since b stands for the y intercept, your parallel equation that passes through (3,-1) is : y = 0x- 1 or just y = -1